OpenStudy (anonymous):

Solve for the variable: -m/4.2 = 5 ? (This is a fraction , sorry for any confusion I'll put the equation in the reply box )

OpenStudy (anonymous):

\[-\frac{ m }{ 4.2 } = 5\]

OpenStudy (anonymous):

so multiply both sides by -4.2

OpenStudy (anonymous):

you multiply by 4.2 on each side and then you divide by negative one on each side

OpenStudy (anonymous):

4.2 x 4.2 = 17.64 4.2 x 5 = 21

OpenStudy (anonymous):

4.2 x 4.2 = 17.64 4.2 x 5 = 21

OpenStudy (anonymous):

dont forget about the minus sign

OpenStudy (anonymous):

why did you write : 4.2 x 4.2 = 17.64 ??

OpenStudy (anonymous):

\[-\frac{ m }{ 4.2 } * (-4.2) = 5 * (-4.2)\]

OpenStudy (anonymous):

I'm not trying to purposely be stupid I'm just horrible at math. I don't get it whatsoever.

OpenStudy (anonymous):

You said multiply both sides?

OpenStudy (anonymous):

do you still need help ? :) and dont worry most people are with practice and time its understandable :D and yes u multiply both by 4.2

OpenStudy (anonymous):

yes.. both sides of the equation we do so to cancel the denominator of the left side and eventually isolate m

OpenStudy (anonymous):

yes I do still need just a bit more help if you could. I'm confused...

OpenStudy (anonymous):

denumerator **

OpenStudy (anonymous):

\[-\frac{ m }{ 4.2 } * (-4.2) = 5 * (-4.2)\]

OpenStudy (anonymous):

do you know how to go from here ?

OpenStudy (anonymous):

look at the left side of the equation .. can you simplify it ?